OpenStudy (anonymous):

whats the value of the expression -2/5 - (3/7) x (7/10)

OpenStudy (inowalst):

\[-\frac{ 2 }{ 5 }-\frac{ 3 }{ 7 }\times\frac{ 7 }{ 10 }\]

OpenStudy (anonymous):

i dont know :(

OpenStudy (inowalst):

Lets do the first two fractions first. Make the denominators the same. \[\frac{ 2 }{ 5 }\times \frac{ 7 }{ 7 }\] \[\frac{ 3 }{ 7 }\times \frac{ 5 }{ 5 }\] Tell me what you get for each after multiplying.

OpenStudy (anonymous):

2/5 x 7/7= 0.4 3/7 x 5/5 = 0.42857142857 ???

OpenStudy (inowalst):

No. \[2\times7=14\]\[5\times7=35\] \[3\times5=15\]\[7\times5=35\] Now: \[\frac{ 14 }{ 35 }\] and \[\frac{ 15 }{ 35 }\] are your new fractions.

OpenStudy (inowalst):

Remember, the first fraction is still a negative. Now that the denominators are the same, subtract 14-15 and keep the denominator the same.

OpenStudy (anonymous):

14-15= -1

OpenStudy (inowalst):

Good! So now your new fraction is \[-\frac{ 1 }{ 35 }\]

OpenStudy (inowalst):

Do you know how to multiply two fractions with different denominators?

OpenStudy (jhannybean):

\[-\frac{ 2 }{ 5 }-\left(\frac{ 3 }{ 7 }\times\frac{ 7 }{ 10 }\right)\] solve the stuff in parenthesis first since you're multiplying it.

OpenStudy (inowalst):

Oh ^ Okay.